OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

I know the slope is 3. and i got y-1=3(x+2)

OpenStudy (anonymous):

(slope-intercept form)

OpenStudy (anonymous):

simply plugged the the x, y numbers and i got 3x-y=-7 as following 3*-2 -1=-7 and 3*1-10=-7

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw :)